OpenStudy (anonymous):

y=2x + 3 & y=4x + 1 find the intersection poiint how do I do this?

OpenStudy (chaise):

y=2x+3 y=4x+1 We can either graph the two points and see where they intersect, (x,y) will be your answer, or you can solve them simultaneously. y=2x+3 y=4x+1 If we take the two equations away from eachother we eliminate the y value. 0 = -2x+2 -2 = -2x x = 1 If x = 1 and y=4x+1 y=4x+1 y=4*1+1 y = 5 The solution is x = 1 and y=5